Problem 6-36 Comparing Cash Flow Streams [LO1]

You’ve just joined the investment banking firm of Dewey, Cheatum, and Howe. They’ve offered you two different salary arrangements. You can have $85,000 per year for the next two years, or you can have $74,000 per year for the next two years, along with a $30,000 signing bonus today. The bonus is paid immediately, and the salary is paid in equal amounts at the end of each month.

  

If the interest rate is 9 percent compounded monthly, what is the value today of each option? (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)
Answer #1

Option 1

PVOrdinary Annuity = C*[(1-(1+i/100)^(-n))/(i/100)]
C = Cash flow per period
i = interest rate
n = number of payments
PV= 7083.3333*((1-(1+ 9/1200)^(-2*12))/(9/1200))
PV = 155048.12

Option 2

PVOrdinary Annuity = C*[(1-(1+i/100)^(-n))/(i/100)]
C = Cash flow per period
i = interest rate
n = number of payments
PV= 6166.6666*((1-(1+ 9/1200)^(-2*12))/(9/1200))
PV = 134983.07

Total = 134983.07+30000=164983.07